OpenStudy (anonymous):

Write the equation of the line passing through (4,7) and parallel to 6x-7y=10

OpenStudy (anonymous):

y = (6/7)x + 25/7 I'm thinking??

OpenStudy (anonymous):

here also you can use y-a=m(y-b) here we know (a,b) through which the line passes..and to find the slope find the slope of the given equation .since two lines are parallel, there slopes are equal

OpenStudy (anonymous):

to find the slope from the given equation you can either convert the equation to the form y=mx+c where m will be the slope...try it

OpenStudy (anonymous):

thx!!

OpenStudy (anonymous):

welcome:)